13. Find the derivative of the power series: f(x) => (-1)(x - 2) - Write the first four terms of f.
Answer #1


۔ 2 - 1 7 ( مامی) 3 (*) = (2-2) + (k )۔ حل (1) = و 2 (2-اس» با (1) = کد تور و ا جرای